{F} A company A lended some day 19006 € to the company B. The interest rate was agreed to be 10 % annually. Later on, the company B paid back to the company A 19483 €. How long was the loan period in calendar days?


1
Expert's answer
2022-02-23T10:18:52-0500

Solution:

Amount, A=19483 €

Principle, P=19006 €

R=10 % annually

Assume we have compounded interest here.

"A=P(1+\dfrac{R}{100})^n
\\ 19483=19006(1+\dfrac{10}{100})^n
\\ 1.025=(1.1)^n
\\ \log 1.025=n\log 1.1
\\ n=0.259"

n is in years here.

"n=0.259\times 365.25\ days=94.6\ days"